# Fluid dynamics, Statistical physics, Theoretical physics

### v2

@@ 7.1 Visualizing Quantized Scalar Shells
Figure 5: The ’ramp-cliff’ structure in the time trace of temperature fluctuations ($\Delta\theta$) in a heated turbulent jet. These asymmetric patterns, characterized by a gradual rise (the ’ramp’) followed by a sharp drop (the ’cliff’), are a key signature of large-scale coherent structures imprinting on small-scale scalar fields. This empirical observation is analogous to the sawtooth profile predicted for the quantized scalar shells in the Euler ensemble (Figure 3) if you reflect this picture across the right-hand side.( Figure adapted from [Sreenivasan(2019b)]
@@
Presumably, these shell structures manifest as the "ramp-cliff" structures with similar sawtooth profiles, observed long ago but never completely understood from the microscopic theory ( see [Sreenivasan(2019b)], (see Figure 5)
@@ 7.2 Confirming Quantum Interference Oscillations
Standard multifractal phenomenology has no mechanism to produce such oscillations. However, remarkably, precisely these types of oscillations have been independently observed in recent, high-precision experiments [Küchler et al.(2023)Küchler, Bewley, and Bodenschatz] on decaying turbulence by the Max Planck Institute group (see Figure 6, Bottom). This data was not published in that paper, but the authors provided it at my request. I shared it with their permission. here.
Figure 7: Left: The turbulent kinetic energy decaying with time in the $4K$ grid simulation with initial Gaussian velocity field distributed according to the K41 spectrum $k^{-5/3}$. After a few hundred time steps, the statistical equilibrium was reached, and energy started decaying in agreement with our $t^{-5/4}$ law. There are other parameters of that stage of decay, with distribution matching our theory, such as the length scale $L(t)\sim\sqrt{t}$ and effective index of the second moment of velocity field which is a nontrivial function of scaling variable $\log(r/\sqrt{t})$ (see Fig. 2). Right: Typical setup of the grid turbulence experiment. The flow enters from the left, passes through the oscillating grid, and creates vorticity. These vortexes interact, exchange energy, and eventually reach statistical equilibrium, with energy constantly dissipated in the bulk while being resupplied by the flow through the grid. In the still Galilean frame, this is a steady process, with energy coming from the left boundary and dissipated in the bulk. The local kinetic energy decays with the distance $z$ from the grid as $z^{-1.25}$, but stays constant with time. In the frame moving with the mean flow velocity, this is decaying turbulence, with energy decaying with time as $(vt)^{-1.25}$.
@@ 8 A New Perspective on Classical Models in Decaying Turbulence
A classic setting is grid turbulence (Fig. 7). The landmark experiments of [Comte-Bellot and Corrsin(1966)] measured a decay exponent $1.25\pm 0.01$, close to $5/4$ and inconsistent with the Kolmogorov–Saffman value $6/5$. More recently, DNS on $4\mathrm{K}$ grids with periodic boundaries and K41-type initial spectra (no forcing) were performed by Sreenivasan and collaborators. The simulations relaxed to a decaying state in which the energy decays as $t^{-5/4}$, the integral scale obeys $L(t)\sim t^{1/2}$, and the effective index of the second moment follows the predicted nonlinear curve (Fig. 7,Left), 2).
@@
  • Analytic foundation. Numerical simulations can test or falsify models but cannot derive the law itself from first principles. The analytic Hodge–dual matrix surface, constructed by the harmonic–map method, exactly solves the fixed–point Yang–Mills loop equation [Migdal(2025a)]. It provides the theoretical explanation for the confining area law rather than an empirical fit, offering a genuine analytic counterpart to the numerical area law of lattice QCD.
@@
With respect to the decaying regime, the classical K41 template is not supported by data at the level often assumed. Kolmogorov’s original constant–dissipation ansatz was a model assumption, and even he (with Obukhov) introduced fluctuations (log–normal) soon after. Modern experiments and DNS [Yakhot and Zakharov(1993), Yakhot and Sreenivasan(2004), Sreenivasan and Yakhot(2021)] show significant, systematic deviations from K41 (and from its log–normal variant). In decaying turbulence, experiments [Panickacheril John et al.(2022a)Panickacheril John, Donzis, and Sreenivasan] deviate even more strongly, despite attempts to fit or trim the data; for example, the energy spectrum departs from $k^{-\frac{5}{3}}$ over many decades, and recent measurements [Küchler et al.(2023)Küchler, Bewley, and Bodenschatz] report clear departures of the "effective index" of $\langle\Delta v^{2}\rangle(r)$ from the K41 value of $2/3$.
@@
Motivated by analogies with critical phenomena, multifractal models posit power laws with anomalous exponents in an inertial range and can organize deviations from K41 in forced turbulence [Parisi and Frisch(1985), Yakhot and Sreenivasan(2004), Sreenivasan and Yakhot(2021)]. In the incompressible turbulence, however, there is no theoretical basis for conformal symmetry: the conservation constraints of a CFT would assign equal dimensions d1=2d-1=2 to divergence-less fields $\boldsymbol{v},\boldsymbol{\omega}$. This relation is incompatible with $\boldsymbol{\omega}=\nabla\times\mathbf{v}$, which would lead to dimensions differing by $1$ rather than being equal. More importantly, the decaying spectra are not straight lines on log–log plots; they bend universally (cf. Fig. 2 (top) of [Küchler et al.(2023)Küchler, Bewley, and Bodenschatz]), indicating that single power laws—whether K41 or multifractal—are insufficient. The loop theory predicts nonlinear, universal functions of $\log k\sqrt{t}$ across several decades, consistent with the observed curvature.
@@ A Generalization of Multifractal Laws
While not conformally invariant, loop space theory provides a microscopic basis for and a generalization of the multifractal paradigm: the multifractal models introduce a continuous spectrum of anomalous exponents phenomenologically, but the loop theory derives, from first principles, an infinite, discrete, and arithmetic spectrum of intermittency exponents. The Mellin transform of the vorticity correlation is found to be meromorphic, yielding the set of scaling exponents shown below:
indexes of velocity correlation
@@
This result fundamentally refines the multifractal idea. The leading exponents in this spectrum, consistent with numerical data, can be viewed as the "ground state" of the scaling law, describing the dominant behavior. The infinite series of subsequent exponents then dictates a hierarchy of calculable, quantum-like corrections. Recent DNS [Sreenivasan(2023)] report slopes consistent with these predictions, in contrast to the classical scaling templates of Kolmogorov or Saffman.
@@
The present framework calls into question not only the K41 scaling laws but also the conventional concept of a "Kolmogorov energy flux." The rapid decay of the energy spectrum predicted by the theory, in accordance with bounds established by Sulem and Frisch [Sulem and Frisch(1975)], corresponds to a vanishing Kolmogorov flux in the turbulent limit. The apparent paradox of finite dissipation in the absence of a spectral flux is resolved by the dissipative anomaly, a mechanism not accounted for in that earlier work.
This mechanism suggests an alternative physical picture to the Richardson cascade. Rather than a scale-by-scale transfer, energy injected at the largest scales (e.g., from boundaries) is dissipated directly on fine-scale, viscous micro-structures. The exact expressions for dissipation and the energy spectrum derived from the loop space approach [Migdal(2024h)] are consistent with the dissipative anomaly and this picture of direct dissipation, but not with a constant-flux energy cascade.
@@
Heisenberg’s $k^{-7}$ proposal for a dissipative subrange, taken up by Chandrasekhar, was a model of its time [Heisenberg(1948), Chandrasekhar(1949)]. As Sreenivasan recounts [Sreenivasan(2019a)], von Neumann already noted in 1949 the lack of experimental support. In our Mellin spectrum, there is no pole between 13/2-13/2 and $8\pm\mathrm{i}t_{n}$, and the "effective index" approaches 7/2-7/2 rather than 7-7. In other words, the nontrivial dynamics persist across what used to be separated as “inertial” and “dissipative” ranges. The resulting universal decaying spectrum spans several decades, and the "effective index" of $\langle\Delta v^{2}\rangle(r)$ varies smoothly from $2$ down to 0, with no plateau at $2/3$. These predictions match DNS and experiments across the entire turbulent range without adjustable dimensionless parameters.
@@
The applicability of results from forced simulations to universal turbulence characteristics requires careful consideration. With finite forcing kept active, the resulting statistical state can depend significantly on the specifics of the forcing scheme (e.g., its spatial structure, correlation time, and the mechanism of energy injection), potentially violating universality (see, e.g., the discussion in [Ishihara et al.(2009)Ishihara, Gotoh, and Kaneda]). In particular, large-wavelength forcing, especially if targeting the potential (irrotational) velocity component, can potentially seed non-universal large–scale structures influenced by long–range correlations $\langle vv\rangle(r)\sim r^{\alpha}\to\infty$. In such cases, the simulated flow may reflect properties of the driving mechanism more than the intrinsic fluid dynamics, akin to altering the taste of soup with the greasy spoon used to stir it.
By contrast, the loop equation isolates the vorticity sector, whose correlations inherently decay with distance and are expected to be less sensitive to boundary conditions and to details of distant forcing. This separation helps identify a universal decaying–turbulence attractor. Crucially, this attractor is directly relevant even to ostensibly steady turbulent flows, such as Feynman’s pipe flow, experimental grid turbulence, and jets. These flows typically exhibit spatial decay downstream from the generation source. By Galilean invariance, in a frame moving with the mean flow velocity $v$, this spatial decay is equivalent to the temporal decay problem addressed by our theory, with downstream distance $z$ corresponding to time $t=z/v$ and the boundary condition at the energy source, such as the inlet of the tube, becoming an initial condition at $t=0$ when a given element of water passes the energy source. The observed spatial energy decay in grid turbulence, $E(z)\sim z^{-5/4}$ [Comte-Bellot and Corrsin(1966)], provides a compelling example, matching the temporal $E(t)\sim t^{-5/4}$ decay predicted by the Euler ensemble. Thus, the study of temporally decaying turbulence offers a first-principles approach to understanding the universal aspects of these physically ubiquitous, spatially developing flows, addressing the long-standing challenges highlighted by Feynman.
@@
In our theory, the loop diffusion equation replaces the Liouville equation, and the fixed-point (decaying) solution that replaces the Gibbs distribution is the Euler ensemble. The analog of the ergodic hypothesis is the conjecture that every distinct state of the Euler ensemble contributes equally to long-time averages, implying that each state is visited with equal frequency over time. We have not proven this conjecture; the closest approach involved obtaining a transcendental equation for the spectrum of the time decay of deviations from the Euler ensemble [Migdal(2023)]. A proof of ergodicity (or convergence to the attractor) would need to demonstrate the stability of the Euler ensemble, i.e., the positivity of the real parts of all the decay indices in its spectrum.
Lacking such a proof, we turn to physical and numerical experiments. These consistently show convergence from initial states with large vorticity towards a solution whose statistical properties closely match the predictions derived from the Euler ensemble, including decay indices and spectra, as discussed in previous sections [Comte-Bellot and Corrsin(1966), Panickacheril John et al.(2022b)Panickacheril John, Donzis, and Sreenivasan, Sreenivasan and Rodhiya, Akash(2026)]. While the stability and ergodic properties of the Euler ensemble remain a mathematical challenge, this empirical agreement provides strong support for its role as the relevant invariant measure for decaying turbulence.
@@ 9 Conclusion and Outlook
Among these avenues, the theory’s formulation of the confining string is of particular interest for quantum chromodynamics. The framework yields an exact solution to the QCD loop equations for non-intersecting loops in the form of a Hodge-dual minimal surface [Migdal(2025c)]. This geometric object, by its very construction, satisfies the loop equations in four-dimensional space, unlike the conventional Euclidean minimal surface, which violates the full set of Yang-Mills loop equations.
The analytic Hodge–dual matrix surface derived by the harmonic–map method [Migdal(2025a)] provides the geometric core of the Yang–Mills fixed–point solution, completing the dual correspondence between turbulence and confinement. Beyond this analytic construction, the holomorphic reduction yields an explicit Confinement Theorem. Define the spectral ratio
$r[C]$ $=\frac{\lambda_{3}+\lambda_{4}}{\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}}\in\Big[\tfrac{1}{2},\,1\Big],$ (60)
where $\lambda_{1}\leq\lambda_{2}\leq\lambda_{3}\leq\lambda_{4}$ are the ordered eigenvalues of the Douglas–Gram tensor
$G_{\mu\nu}[C]$ $=\sum_{n>0}n\,\hat{C}_{n,\mu}\,\hat{C}_{n,\nu},\qquad\hat{C}_{n,\mu}=\tfrac{1}{2\pi}\!\int_{0}^{2\pi}\!C_{\mu}(\theta)e^{-in\theta}\,d\theta.$ (61)
From the analytic solution, the Hodge–dual minimal area of the surface and the Euclidean minimal area of the same boundary loop are expressed as
$S_{\chi}[C]$ $=2\sqrt{2}\,(\lambda_{3}+\lambda_{4}),$ $S_{\mathrm{E}}[C]$ $=\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4},$ (62)
so that their ratio directly yields the spectral ratio $r[C]$. Hence the effective string tension,
$\sigma_{\mathrm{eff}}[C]$ $=\frac{S_{\chi}[C]}{S_{\mathrm{E}}[C]}=2\sqrt{2}\,r[C],$ (63)
obeys the universal bounds
$\sqrt{2}\ \leq\ \sigma_{\mathrm{eff}}[C]\ \leq\ 2\sqrt{2},$ (64)
for all loops. The numerical value of the lower bound is simple, but its existence—independent of loop size and shape—is highly nontrivial and constitutes the confinement statement in this framework. All residual shape dependence of the string tension is encoded by the spectral ratio $r[C]$: planar loops (rank–2 $G$) saturate the upper bound $r[C]=1$, while isotropic spectra saturate the lower bound $r[C]=\tfrac{1}{2}$. This explicit formula for the shape dependence of the quark string tension opens the way to computing the rising Regge trajectories in QCD from first principles.

### v3

@@ 7.1 Visualizing Quantized Scalar Shells
Figure 6: The ’ramp-cliff’ structure in the time trace of temperature fluctuations ($\Delta\theta$) in a heated turbulent jet. These asymmetric patterns, characterized by a gradual rise (the ’ramp’) followed by a sharp drop (the ’cliff’), are a key signature of large-scale coherent structures imprinting on small-scale scalar fields. This empirical observation is analogous to the sawtooth profile predicted for the quantized scalar shells in the Euler ensemble (Figure 4) if you reflect this picture across the right-hand side.( Figure adapted from [sreenivasan2018turbulent]
@@
Presumably, these shell structures manifest as the "ramp-cliff" structures with similar sawtooth profiles, observed long ago but never completely understood from the microscopic theory ( see [sreenivasan2018turbulent], (see Figure 6)
@@ 7.2 Confirming Quantum Interference Oscillations
Standard multifractal phenomenology has no mechanism to produce such oscillations. However, remarkably, precisely these types of oscillations have been independently observed in recent, high-precision experiments [GregXi2] on decaying turbulence by the Max Planck Institute group (see Figure 7, Bottom). This data was not published in that paper, but the authors provided it at my request. I shared it with their permission. here.
Figure 8: Left: The turbulent kinetic energy decaying with time in the $4096^{3}$ grid simulation with initial Gaussian velocity field distributed according to the K41 spectrum $k^{-5/3}$. After a few hundred time steps, the statistical equilibrium was reached, and energy started decaying in agreement with our $t^{-5/4}$ law. There are other parameters of that stage of decay, with distribution matching our theory, such as the length scale $L(t)\sim\sqrt{t}$ and effective index of the second moment of velocity field which is a nontrivial function of scaling variable $\log(r/\sqrt{t})$ (see Fig. 2). Right: Typical setup of the grid turbulence experiment. The flow enters from the left, passes through the oscillating grid, and creates vorticity. These vortexes interact, exchange energy, and eventually reach statistical equilibrium, with energy constantly dissipated in the bulk while being resupplied by the flow through the grid. In the still Galilean frame, this is a steady process, with energy coming from the left boundary and dissipated in the bulk. The local kinetic energy decays with the distance $z$ from the grid as $z^{-1.25}$, but stays constant with time. In the frame moving with the mean flow velocity, this is decaying turbulence, with energy decaying with time as $(vt)^{-1.25}$.
@@ 8 A New Perspective on Classical Models in Decaying Turbulence
A classic setting is grid turbulence (Fig. 8). The landmark experiments of [GridTurbulence_1966] measured a decay exponent $1.25\pm 0.01$, close to $5/4$ and inconsistent with the Kolmogorov–Saffman value $6/5$. More recently, DNS on $4\mathrm{K}$ grids with periodic boundaries and K41-type initial spectra (no forcing) were performed by Sreenivasan and collaborators. The simulations relaxed to a decaying state in which the energy decays as $t^{-5/4}$, the integral scale obeys $L(t)\sim t^{1/2}$, and the effective index of the second moment follows the predicted nonlinear curve (Fig. 8,Left), 2).
@@
  • Analytic foundation. Numerical simulations can test or falsify models but cannot derive the law itself from first principles. The analytic Hodge–dual matrix surface, constructed by the harmonic–map method, exactly solves the fixed–point Yang–Mills loop equation [migdal2025HodgeDualMatrix]. It provides the theoretical explanation for the confining area law rather than an empirical fit, offering a genuine analytic counterpart to the numerical area law of lattice QCD.
@@
With respect to the decaying regime, the classical K41 template is not supported by data at the level often assumed. Kolmogorov’s original constant–dissipation ansatz was a model assumption, and even he (with Obukhov) introduced fluctuations (log–normal) soon after. Modern experiments and DNS [YZ93, YS04, SY21] show significant, systematic deviations from K41 (and from its log–normal variant). In decaying turbulence, experiments [SreeniDecaying] deviate even more strongly, despite attempts to fit or trim the data; for example, the energy spectrum departs from $k^{-\frac{5}{3}}$ over many decades, and recent measurements [GregXi2] report clear departures of the "effective index" of $\langle\Delta v^{2}\rangle(r)$ from the K41 value of $2/3$.
@@
Motivated by analogies with critical phenomena, multifractal models posit power laws with anomalous exponents in an inertial range and can organize deviations from K41 in forced turbulence [FP85, YS04, SY21]. In the incompressible turbulence, however, there is no theoretical basis for conformal symmetry: the conservation constraints of a CFT would assign equal dimensions d1=2d-1=2 to divergence-less fields $\boldsymbol{v},\boldsymbol{\omega}$. This relation is incompatible with $\boldsymbol{\omega}=\nabla\times\mathbf{v}$, which would lead to dimensions differing by $1$ rather than being equal. More importantly, the decaying spectra are not straight lines on log–log plots; they bend universally (cf. Fig. 2 (top) of [GregXi2]), indicating that single power laws—whether K41 or multifractal—are insufficient. The loop theory predicts nonlinear, universal functions of $\log k\sqrt{t}$ across several decades, consistent with the observed curvature.
@@ A Generalization of Multifractal Laws
While not conformally invariant, loop space theory provides a microscopic basis for and a generalization of the multifractal paradigm: the multifractal models introduce a continuous spectrum of anomalous exponents phenomenologically, but the loop theory derives, from first principles, an infinite, discrete, and arithmetic spectrum of intermittency exponents.
$\text{CFT: }\vec{v}(0)\cdot\vec{v}(r)\sim\sum_{n}O_{n}|r|^{p_{n}};\;p_{n}\stackrel{{\scriptstyle?}}{{=}}2\Delta_{v}-\Delta_{n};$
The Mellin transform of the velocity correlation is found to be meromorphic, yielding the set of scaling exponents shown below:
indexes $p_{n}$ of velocity correlation
@@
This result fundamentally refines the multifractal idea. The leading exponents in this spectrum, consistent with numerical data, can be viewed as the "ground state" of the scaling law, describing the dominant behavior. The infinite series of subsequent exponents then dictates a hierarchy of calculable, quantum-like corrections. Recent DNS [DecayingTalk] report slopes consistent with these predictions, in contrast to the classical scaling templates of Kolmogorov or Saffman.
@@
The present framework calls into question not only the K41 scaling laws but also the conventional concept of a "Kolmogorov energy flux." The rapid decay of the energy spectrum predicted by the theory, in accordance with bounds established by Sulem and Frisch [Sulem_Frisch_1975], corresponds to a vanishing Kolmogorov flux in the turbulent limit. The apparent paradox of finite dissipation in the absence of a spectral flux is resolved by the dissipative anomaly, a mechanism not accounted for in that earlier work.
This mechanism suggests an alternative physical picture to the Richardson cascade. Rather than a scale-by-scale transfer, energy injected at the largest scales (e.g., from boundaries) is dissipated directly on fine-scale, viscous micro-structures. The exact expressions for dissipation and the energy spectrum derived from the loop space approach [migdal2024quantum] are consistent with the dissipative anomaly and this picture of direct dissipation, but not with a constant-flux energy cascade.
@@
Heisenberg’s $k^{-7}$ proposal for a dissipative subrange, taken up by Chandrasekhar, was a model of its time [Heisenberg1948PRSA, Chandrasekhar1949PRSA]. As Sreenivasan recounts [SreeniChandra], von Neumann already noted in 1949 the lack of experimental support. In our Mellin spectrum, there is no pole between 13/2-13/2 and $8\pm\mathrm{i}t_{n}$, and the "effective index" approaches 7/2-7/2 rather than 7-7. In other words, the nontrivial dynamics persist across what used to be separated as “inertial” and “dissipative” ranges. The resulting universal decaying spectrum spans several decades, and the "effective index" of $\langle\Delta v^{2}\rangle(r)$ varies smoothly from $2$ down to 0, with no plateau at $2/3$. These predictions match DNS and experiments across the entire turbulent range without adjustable dimensionless parameters.
@@
The applicability of results from forced simulations to universal turbulence characteristics requires careful consideration. With finite forcing kept active, the resulting statistical state can depend significantly on the specifics of the forcing scheme (e.g., its spatial structure, correlation time, and the mechanism of energy injection), potentially violating universality (see, e.g., the discussion in [Ishihara2009]). In particular, large-wavelength forcing, especially if targeting the potential (irrotational) velocity component, can potentially seed non-universal large–scale structures influenced by long–range correlations $\langle vv\rangle(r)\sim r^{\alpha}\to\infty$. In such cases, the simulated flow may reflect properties of the driving mechanism more than the intrinsic fluid dynamics, akin to altering the taste of soup with the greasy spoon used to stir it.
By contrast, the loop equation isolates the vorticity sector, whose correlations inherently decay with distance and are expected to be less sensitive to boundary conditions and to details of distant forcing. This separation helps identify a universal decaying–turbulence attractor. Crucially, this attractor is directly relevant even to ostensibly steady turbulent flows, such as Feynman’s pipe flow, experimental grid turbulence, and jets. These flows typically exhibit spatial decay downstream from the generation source. By Galilean invariance, in a frame moving with the mean flow velocity $v$, this spatial decay is equivalent to the temporal decay problem addressed by our theory, with downstream distance $z$ corresponding to time $t=z/v$ and the boundary condition at the energy source, such as the inlet of the tube, becoming an initial condition at $t=0$ when a given element of water passes the energy source. The observed spatial energy decay in grid turbulence, $E(z)\sim z^{-5/4}$ [GridTurbulence_1966], provides a compelling example, matching the temporal $E(t)\sim t^{-5/4}$ decay predicted by the Euler ensemble. Thus, the study of temporally decaying turbulence offers a first-principles approach to understanding the universal aspects of these physically ubiquitous, spatially developing flows, addressing the long-standing challenges highlighted by Feynman.
@@
In our theory, the loop diffusion equation replaces the Liouville equation, and the fixed-point (decaying) solution that replaces the Gibbs distribution is the Euler ensemble. The analog of the ergodic hypothesis is the conjecture that every distinct state of the Euler ensemble contributes equally to long-time averages, implying that each state is visited with equal frequency over time. We have not proven this conjecture; the closest approach involved obtaining a transcendental equation for the spectrum of the time decay of deviations from the Euler ensemble [migdal2023exact]. A proof of ergodicity (or convergence to the attractor) would need to demonstrate the stability of the Euler ensemble, i.e., the positivity of the real parts of all the decay indices in its spectrum.
Lacking such a proof, we turn to physical and numerical experiments. These consistently show convergence from initial states with large vorticity towards a solution whose statistical properties closely match the predictions derived from the Euler ensemble, including decay indices and spectra, as discussed in previous sections [GridTurbulence_1966, TurbulentDecaySreeni22, SreeniAkash2025]. While the stability and ergodic properties of the Euler ensemble remain a mathematical challenge, this empirical agreement provides strong support for its role as the relevant invariant measure for decaying turbulence.
Remark 12.
One of the thoughtful readers of this manuscript asked the following question:
Finally, considering the quantum and linear nature of the diffusion-loop equation, does that author have any comments on the possibility of utilizing quantum computers to study these systems? …
This question resonates deeply with our own experience. During the long struggle with the Navier-Stokes loop equation—before the analytic simplicity and beauty of the Euler ensemble emerged—we explored various numerical avenues. Even now, after two years of successfully investigating the properties of the Euler ensemble in infinite space, the necessity for a powerful numerical method remains acute for finite geometries.
Quantum computing indeed offers a unique opportunity here. The chance lies not in verifying the Euler ensemble (which is exact for the infinite case), but in solving the loop diffusion (Schrödinger) equation in a finite geometry by mapping it onto a quantum mechanical system of qubits. This approach appears far more efficient than the existing proposals for mapping the nonlinear Navier-Stokes equation onto a grid of qubits. In the case of the loop equation, we possess a linear Hamiltonian structure (albeit non-Hermitian) from the outset. The qubits would live on a loop, moving in a physical space, which looks like a huge reduction of dimensionality.
We have made several passes towards experts in quantum simulations, only to find that, for the moment, such simulations remain at the level of science fiction. However, a breakthrough in quantum hardware could occur at any time. Therefore, we believe it is vital to keep this unique path open: solving for large-scale turbulent flows around physical bodies by simulating the loop equation on a future quantum computer.
@@ 9 Conclusion and Outlook
Among these avenues, the theory’s formulation of the confining string is of particular interest for quantum chromodynamics. The framework yields an exact solution to the QCD loop equations in the form of a Hodge-dual minimal surface [migdal2025YangMills]. This geometric object, by its very construction, satisfies the loop equations in four-dimensional space, unlike the conventional Euclidean minimal surface, which violates the full set of Yang-Mills loop equations.
The analytic Hodge–dual matrix surface derived by the harmonic–map method [migdal2025HodgeDualMatrix] provides the geometric core of the Yang–Mills fixed–point solution, completing the dual correspondence between turbulence and confinement.
This explicit formula for the shape dependence of the quark string evolution opens the way to computing the rising Regge trajectories in QCD from first principles.